### v6

@@ 5 Conclusions and Future Works
This paper investigates bilevel optimization without the typical lower-level strong convexity assumption. We have shown that finding points with small hyper-gradients is computationally hard for nonconvex-convex bilevel problems, but easy for nonconvex-PL bilevel problems where simple first-order algorithms can achieve fast convergence rates. Finally, we discuss possible future directions.
Problem Complexity.
Assumption 4.1d suppose $f$ has $\rho_{f}$-Lipschitz Hessians, which comes from Lemma 4. Such an assumption is not required by nonconvex-strongly-convex problems (Chen et al., 2023). We wonder if it is possible to remove this assumption in our setting. Also, it is also interesting to study lower bounds for this problem (Kwon et al., 2024a).
More Lower-Level Problems
It is also interesting to study bilevel problems with lower-level functions beyond the PL condition. One possible direction is to consider the Kurdyka-Łojasiewicz (KL) condition (Fatkhullin et al., 2022), which is more general than the PL condition and it holds for any semialgebraic functions. Recent works have shown non-asymptotic convergence for nonconvex-KL minimax optimization problems (Li et al., 2022; Zheng et al., 2023), but the nonconvex-KL bilevel optimization problem remains challenging. Besides these conditions, we hope future works to exploit the problem structure that arises from the applications of bilevel optimization and identify other tractable conditions that can break the limit of our hardness result.
@@ Appendix A Related Works
There have been many works studying bilevel optimization. Here we only review related works focusing on non-asymptotic analysis.
@@
and applies gradient descent on the penalty function. Several studies have demonstrated two notable advantages of this method over AID and ITD: Firstly, it only requires gradient oracles which are much cheaper than Hessian-vector product oracles (Liu et al., 2022a). Secondly, it performs well empirically even in cases where the lower-level function is nonconvex (Liu et al., 2021a; Shen and Chen, 2023). By establishing the equivalence between the KKT (Karush–Kuhn–Tucker) point of Problem (10) and the stationary point of Problem (2), Kwon et al. (2023) showed that F2BA can find an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls when the lower-level function is strongly convex. Chen et al. (2023) further improved the rate to $\tilde{\mathcal{O}}(\epsilon^{-2})$ by a more careful landscape analysis of the penalty function. Kwon et al. (2024b); Yao et al. (2024) proposed proximal variants of F2BA by replacing the lower-level value function with its Moreau envelope to tackle lower-level constraints. Kwon et al. (2024b) showed a $\tilde{\mathcal{O}}(\epsilon^{-3})$ complexity when the penalty function satisfies the Prox-EB condition (which is equivalent to the PL condition in the unconstrained case as Proposition 1).
@@
Xiao et al. (2023) showed GALET can converge to a KKT point of Problem (11) with a rate of $\tilde{\mathcal{O}}(\epsilon^{-2})$. In Appendix B, we show that this also implies the same convergence rate to the stationary point of $\varphi(x)$ under the assumptions of Kwon et al. (2024b).
Huang (2023) also considers nonconvex-PL bilevel problems. Our work differs from (Huang, 2023) in the following aspects. First, Assumption 2 in (Huang, 2023) supposes that $\nabla^{2}g(x,y)$ is non-singular at the minimizer of g, while we do not require such an assumption. Secondly, our algorithm only requires first-order oracles, while (Huang, 2023) requires second-order oracles. Thirdly, our algorithm has only ${\mathcal{O}}(d)$ complexity at each step, where $d=\max\{d_{x},d_{y}\}$ is the dimension of the problem. In contrast, each iteration of the algorithm in (Huang, 2023) requires a ${\mathcal{O}}(d^{3})$ complexity for computing the SVD decomposition of the Hessian matrix.
@@
Remark B.2
Xiao et al. (2023) additionally assumes the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero when solving Equation (12). However, the complexity to solve Equation (12) only depends on the singular value of $\nabla_{yy}^{2}g(x,y^{*})$, which can be bound below by our Lemma 6. Since $g(x,y)$ is PL in $y$, by Lemma 8 gradient descent on $g$ can make ${\rm dist}(y^{k},Y^{*}(x))$ converges linearly. So one can let $y^{K}$ sufficiently close to some point in $Y^{*}(x)$ with logarithmic large $K$. Then by Weyl’s inequality the singular value of $\nabla^{2}_{yy}g(x,y^{K})$ is also bounded below. Therefore, such an additional assumption in Xiao et al. (2023) seems to be redundant.
Appendix C Discussion on Kwon et al. (2024b)
Different from our Assumption 4.1a, Kwon et al. (2024b) uses the following proximal error bound (Prox-EB) assumption.
@@ Appendix F Proof of Theorem 4
It follows the generalized Danskin’s theorem proved in Shen and Chen (2023). See also Lemma A.2 in Kwon et al. (2024b).
@@
The proof follows Theorem 3.8 Kwon et al. (2024b). The only difference is that Theorem 3.8 Kwon et al. (2024b) states $\|\nabla\varphi_{\sigma}(x)-\varphi(x)\|={\mathcal{O}}(\sigma\ell\kappa^{5})$. But the additional $\kappa^{2}$ dependency comes from the perturbation in the multiplier ${\rm d}\lambda/{\rm d}\sigma\asymp\kappa^{2}$. Since we only consider the unconstrained case, there is no need to consider the effect of ${\rm d}\lambda/{\rm d}\sigma$, and we can improve the bound to ${\mathcal{O}}(\sigma\ell\kappa^{3})$.
We also recall some technical lemmas from Kwon et al. (2024b).
@@
Finally, we plug in the explicit form of $\dfrac{\partial^{2}}{\partial x\partial\sigma}l(x,\sigma)$ by Theorem 3.2 (Kwon et al., 2024b).
@@ Appendix G Proof of Theorem 7
${\mathbb{E}}\left[{\rm dist}^{2}(x_{T},X^{*})\right]\leq\left(1-\frac{\alpha}{\beta}\right)^{T}\frac{\beta}{\alpha}{\rm dist}^{2}(x_{0},X^{*})+\frac{M^{2}}{\alpha\beta B}.$
@@
${\mathbb{E}}\left[h(x_{t+1})-h^{*}\right]\leq\left(1-\frac{\alpha}{\beta}\right)^{T}\left(h(x_{0})-h^{*}\right)+\frac{M^{2}}{2\beta B}.$
@@
$\sum_{t=0}^{T-1}\delta_{t}\leq 2\delta_{0}+\frac{32L_{g}^{2}}{\mu^{2}}\sum_{t=0}^{T-1}\|x_{t+1}-x_{t}\|^{2}+\frac{\sigma^{2}\epsilon^{2}}{2L_{g}^{2}},$

### v8

@@ 5 Conclusions and Future Works
This paper investigates bilevel optimization without the typical lower-level strong convexity assumption. We have shown that finding points with small hyper-gradients is computationally hard for nonconvex-convex bilevel problems, but easy for nonconvex-PL bilevel problems where simple first-order algorithms can achieve fast convergence rates.
It will be interesting to study bilevel problems with lower-level functions beyond the PL condition in the future. One possible direction is to consider the Kurdyka-Łojasiewicz (KL) condition (Fatkhullin et al., 2022), which is more general than the PL condition and it holds for any semialgebraic functions. Recent works have shown non-asymptotic convergence for nonconvex-KL minimax optimization problems (Li et al., 2022; Zheng et al., 2023), but the nonconvex-KL bilevel optimization problem remains challenging. Besides these conditions, we hope future works to exploit the problem structure that arises from the applications of bilevel optimization and identify other tractable conditions that can break the limit of our hardness result.
@@ Appendix A Related Works
and applies gradient descent on the penalty function. Several studies have demonstrated two notable advantages of this method over AID and ITD: Firstly, it only requires gradient oracles which are much cheaper than Hessian-vector product oracles (Liu et al., 2022a). Secondly, it performs well empirically even in cases where the lower-level function is nonconvex (Liu et al., 2021a; Shen and Chen, 2023). By establishing the equivalence between the KKT (Karush–Kuhn–Tucker) point of Problem (10) and the stationary point of Problem (2), Kwon et al. (2023) showed that F2BA can find an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls when the lower-level function is strongly convex. Chen et al. (2023) further improved the rate to $\tilde{\mathcal{O}}(\epsilon^{-2})$ by a more careful landscape analysis of the penalty function. Kwon et al. (2024); Yao et al. (2024) proposed proximal variants of F2BA by replacing the lower-level value function with its Moreau envelope to tackle lower-level constraints. Kwon et al. (2024) showed a $\tilde{\mathcal{O}}(\epsilon^{-3})$ complexity when the penalty function satisfies the Prox-EB condition (which is equivalent to the PL condition in the unconstrained case as Proposition 1).
@@
Xiao et al. (2023) showed GALET can converge to a KKT point of Problem (11) with a rate of $\tilde{\mathcal{O}}(\epsilon^{-2})$. In Appendix B, we show that this also implies the same convergence rate to the stationary point of $\varphi(x)$ under the assumptions of Kwon et al. (2024).
Huang (2023, 2024) also considers nonconvex-PL bilevel problems. Our work differs from these works in the following aspects. First, Assumption 2 in (Huang, 2023) $g(x,y)$ has a unique minimizer $y^{*}(x)$ and $\nabla_{yy}^{2}g(x,y^{*}(x))$ is non-singular, which is much easier than our setting as we have discussed in Remark 4.2. Secondly, our algorithm only requires first-order oracles, while (Huang, 2023) requires second-order oracles. Thirdly, our algorithm has only ${\mathcal{O}}(d)$ complexity at each step, where $d=\max\{d_{x},d_{y}\}$ is the dimension of the problem. In contrast, each iteration of the algorithm in (Huang, 2023) requires a ${\mathcal{O}}(d^{3})$ complexity for computing the SVD decomposition of the Hessian matrix. Although Huang (2024) claimed a projection operator can remove the expensive SVD decomposition, the claim seems to be incorrect. Huang (2024) defined projector ${\mathcal{M}}(Hv)$ such that ${\mathcal{M}}(Hv)={\mathcal{S}}_{[\mu,L_{g}]}(H){\mathcal{P}}_{r}(v)$, where ${\mathcal{S}}_{[\mu,L_{g}]}(H)$ projects all the singular-values of $H$ into the interval $[\mu,L_{g}]$ and ${\mathcal{P}}_{r}(v)$ projects vector $v$ onto the set $\{v\in{\mathbb{R}}^{d_{y}}:\|v\|\leq r\}$. Consider $H=\begin{pmatrix}2&0\\ 0&1\end{pmatrix}$, $v=\begin{pmatrix}1\\ 1\end{pmatrix}$, $L=\frac{3}{2}$, $\mu=\frac{1}{2}$, $r_{v}=\sqrt{2}$ then $Hv=\begin{pmatrix}2\\ 1\end{pmatrix}$, but ${\mathcal{M}}(Hv)=\begin{pmatrix}\frac{3}{2}\\ 1\end{pmatrix}$ by definition. Hence, ${\mathcal{M}}(Hv)\neq{\mathcal{P}}_{r^{\prime}}(Hv)$ for any $r^{\prime}$. In fact, it seems that no deterministic function can implement ${\mathcal{M}}(Hv)$. Let $v^{\prime}=\begin{pmatrix}2\\ 1\end{pmatrix}$, $H^{\prime}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. Then ${\mathcal{M}}(H^{\prime}v^{\prime})=\frac{\sqrt{10}}{5}\begin{pmatrix}2\\ 1\end{pmatrix}$ by definition. We have ${\mathcal{M}}(H^{\prime}v^{\prime})\neq{\mathcal{M}}(Hv)$, but $H^{\prime}v^{\prime}=Hv$. Therefore, it seems that the algorithms in (Huang, 2023, 2024) actually both require very expensive Hessian oracles with ${\mathcal{O}}(d^{3})$ running time each iteration, which is different from previous works (and our work) that only use Hessian-vector product oracles or gradient oracles with ${\mathcal{O}}(d)$ running time per iteration.
@@
Appendix C Discussion on Kwon et al. (2024)
Different from our Assumption 4.1a, Kwon et al. (2024) uses the following proximal error bound (Prox-EB) assumption.
@@ Appendix F Proof of Theorem 4
It follows the generalized Danskin’s theorem proved in Shen and Chen (2023). See also Lemma A.2 in Kwon et al. (2024).
@@
The proof follows Theorem 3.8 Kwon et al. (2024). The only difference is that Theorem 3.8 Kwon et al. (2024) states $\|\nabla\varphi_{\sigma}(x)-\varphi(x)\|={\mathcal{O}}(\sigma\ell\kappa^{5})$. But the additional $\kappa^{2}$ dependency comes from the perturbation in the multiplier ${\rm d}\lambda/{\rm d}\sigma\asymp\kappa^{2}$. Since we only consider the unconstrained case, there is no need to consider the effect of ${\rm d}\lambda/{\rm d}\sigma$, and we can improve the bound to ${\mathcal{O}}(\sigma\ell\kappa^{3})$.
We also recall some technical lemmas from Kwon et al. (2024).
@@
Finally, we plug in the explicit form of $\dfrac{\partial^{2}}{\partial x\partial\sigma}l(x,\sigma)$ by Theorem 3.2 (Kwon et al., 2024).
@@ Appendix G Proof of Theorem 7
${\mathbb{E}}\left[{\rm dist}^{2}(x_{T},X^{*})\right]\leq\left(1-\frac{\alpha}{\beta}\right)^{T}\frac{\beta}{\alpha}{\rm dist}^{2}(x_{0},X^{*})+\frac{M^{2}}{\alpha^{2}B}.$
@@
${\mathbb{E}}\left[h(x_{t+1})-h^{*}\right]\leq\left(1-\frac{\alpha}{\beta}\right)^{T}\left(h(x_{0})-h^{*}\right)+\frac{M^{2}}{2\alpha B}.$
@@
$\sum_{t=0}^{T-1}\delta_{t}\leq 2\delta_{0}+\frac{32L_{g}^{2}}{\mu^{2}}\sum_{t=0}^{T-1}\|x_{t+1}-x_{t}\|^{2}+{\mathcal{O}}\left(\frac{\sigma^{2}\epsilon^{2}}{L_{g}^{2}}\right),$